OpenStudy (anonymous):

abcd is a rectangle. find the length of each diagonal.ac=2(5a+1) bd=2(a+1).

OpenStudy (angela210793):

Is there anything else given

OpenStudy (anonymous):

no.

OpenStudy (angela210793):

I'm sorry....nothing's on my mind right now...ask Amistre he/she may know how to solve it

OpenStudy (anonymous):

distribute & add/ sub. like terms from there so: 2(5a+1)=2(a+1) 10a+5=2a+2 (10a-2a)+5=(2a-2a)+2 8a+5=2 8a+(5-5)=(2-5) 8a=-3 a=-3/8

OpenStudy (anonymous):

In a rectangle both diagonal are of equal length. so 2(5a+1)=2(a+1) Opening brackets and expanding, we get 10a+2=2a+2 Separating terms with variable on one side, we get 10a-2a = 2-2 8a = 0 a = 0/8 = 0 so ac = 2(5a+1) = 2(5*0 + 1) = 2(1) = 2 bd = 2(a+1) = 2(0 +1) = 2(1) = 2

OpenStudy (anonymous):

LadiiTreble made a mistake in calculations....

OpenStudy (anonymous):

oh i did woops!

OpenStudy (anonymous):

it happens to best of us, i hv made my own bloopers many a time.....

OpenStudy (anonymous):

Don't worry, be happy :D